Uranus's moon Titania has a radius of 7.89 ✕ 105 m and a measured gravitational field of 0.377 m/s2. What is its mass (in kg)?
Answer #1

Gravitational field of a object is given by:

g = GM/R^2

G = gravitational constant = 6.67*10^-11

M = Mass of object (Titania) = ?

R = radius of Titania = 7.89*10^5 m

So,

M = g*R^2/G

M = 0.377*(7.89*10^5)^2/(6.67*10^-11)

M = 3.52*10^21 kg = Mass of Titania

Please Upvote.